OpenStudy (anonymous):

Find the limit limit as x approaches negative infinity for all of (square root of 4x^6-x)/x^3+5

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ \sqrt{4x ^{6} - x} }{ x^3 + 5 }\] did i write this correctly ?

OpenStudy (anonymous):

er, that should be -inf

OpenStudy (anonymous):

\[\lim_{x \rightarrow- \infty} \frac{ \sqrt{4x ^{6} - x} }{ x^3 + 5 }\]

OpenStudy (anonymous):

\[-\infty \] is the right way but other than that its right

OpenStudy (anonymous):

2 isn't right thanks though

OpenStudy (anonymous):

after several L'Hopital's i got -2